OpenStudy (anonymous):

Which complex number is farthest from the origin in the complex plane? a)6i b)3-2i c)6+i d)-4-7i

OpenStudy (anonymous):

d, by far

OpenStudy (anonymous):

Thankyou!:)

OpenStudy (ash2326):

find the absolute value of each, the highest is the farthest from origin a) 6 b)root(3^2+2^2)=root(13) c)root(6^2+1^2)=root(37) d)root(4^2+7^2)=root(65) so option d